1. Your client presents with preeclampsia. You have on hand Magnesium Sulfate 40 grams in 500 ml of Lactated Ringers. The order is to infuse a maintenance dose of Magnesium Sulfate at 2 grams / hour. How many ml / hour will you run the IV pump?
Answer #1

Ans) 40 gms MgSO4 = 500 ml of Lactated Ringers

2 gms = x

x= 500 × 2/40 = 1000/40 = 25ml/hr

The IV pump to be set at 25 ml/hr.